OpenStudy (anonymous):

Solve for x. Round your answer to two decimal places. Show your work for full credit. A right triangle is shown with the hypotenuse labeled x. One angle has a measure of thirty two degrees, and the opposite leg has a length of twenty.

OpenStudy (anonymous):

heres the picture of the triangle

OpenStudy (anonymous):

\[\sin(\theta) = \frac{ opposite }{ hypotenuse}\]

OpenStudy (anonymous):

\[\sin(32) = \frac{ 20 }{ x }\]

OpenStudy (anonymous):

Actually, I believe x = 37.74

OpenStudy (anonymous):

thank you!! I think youre right @ComputerNerd thats what I got

OpenStudy (anonymous):

37.74 is correct.